Solve each of the following systems:
2x+3y=12 (1)
4x-3y=6 (2)
(1) * 2
4x+6y=24 (3)
(3)-(2)
-9y = -18
Therefore y = 2 x = 3
Use the same principles to solve the other linear equations
y=3x+1
2x-3y=4
3x-2y=7
x=4y-11

Answer by JimboP1977(308)

2x+3y=12 (1)
4x-3y=6 (2)
(1) * 2
4x+6y=24 (3)
(3)-(2)
-9y = -18
Therefore y = 2 x = 3
Use the same principles to solve the other linear equations